OpenStudy (anonymous):

2x + 4 = –3x + 14 help me solve this please????

OpenStudy (bahrom7893):

Add 3x to both sides: 2x + 4 + 3x = -3x + 14 + 3x 2x + 3x + 4 = -3x + 3x + 14 5x + 4 = 0 + 14 5x + 4 = 14

OpenStudy (bahrom7893):

Subtract 4 from both sides: 5x + 4 - 4 = 14 - 4 5x + 0 = 10 5x = 10

OpenStudy (bahrom7893):

what times 5 equals ten?

OpenStudy (anonymous):

2

OpenStudy (bahrom7893):

that's your answer.